The water tank has a maximum capacity of 84 gallons if the tank was 2/7 full how many gallons did it have
Answer by Maths68(1474)

2/7 of 84 = (2/7)(84)
2/7 of 84 = (2)(12)
2/7 of 84 = 24
It has 24 gallons